OpenStudy (anonymous):

Simplify: z3 - 4z /z2 - 4z + 4

OpenStudy (anonymous):

That us supposed to be z^3, and z^2 sorry for the typo

OpenStudy (tkhunny):

Do you remember the Order of Operations?

OpenStudy (anonymous):

pemdas

OpenStudy (tkhunny):

Right. Remembering this, you need to see that you just wrote \(z^{3} - \dfrac{4z}{z^{2}} - 4z + 4\). Do you see that?

OpenStudy (anonymous):

no thats not how it is wrote it is z^3-4z (numerator) / z^2-4z+4 (denominator)

OpenStudy (tkhunny):

You are not seeing it. What you wrote is what I displayed. You can simply add parentheses to make it more clear. (z3 - 4z) /(z2 - 4z + 4) In any case, you need to pull out your Factoring Genius Hat! I'll do the numerator. \(\dfrac{z^{3} - 4z}{z^{2} - 4z + 4} = \dfrac{z(z^{2} - 4)}{z^{2} - 4z + 4} = \dfrac{z(z+2)(z-2)}{z^{2} - 4z + 4}\) You do the denominator.

OpenStudy (anonymous):

would the answer come out to be z(z+2)/z-2 ?

OpenStudy (anonymous):

tkhunny?

OpenStudy (tkhunny):

Except for the missing parentheses in the denominator, again. Can't convince you, can I?